# Reconstructing Hand-Held Objects in 3D from Images and Videos

### v3

@@ 5 Retrieval-Augmented Reconstruction
Like any other deep learning model, how well MCC-HO is expected to perform is related to the amount of training data available, e.g., the size of available hand-object datasets. These are currently all quite small, orders of magnitude smaller than the “big data” on which computer vision foundation models have been trained. How might we find a way to exploit these other models to our end? Specifically, recognition of objects is very well developed [51], and large collections of 3D object models [10, 11] exist and have facilitated advancements in 3D generative modeling [45, 57, 7, 43]. Thus, in order to further improve upon network-inferred object geometry, we exploit these “big data” models by using GPT-4(V) [51] and a text-to-3D generative model [45] to automatically detect and “retrieve” a 3D object corresponding to the hand-held object in an image. We call this technique Retrieval-Augmented Reconstruction (RAR), somewhat in analogy to Retrieval-Augmented Generation in natural language processing [35].
Given a representative frame of the video, we prompt GPT-4(V) to provide a detailed description of the object the hand is holding. This text caption is passed to Genie, a text-to-3D model from Luma [45], to retrieve realistic 3D object geometry and appearance (Genie is also used in [36]). Since Genie outputs geometry normalized to fit inside a unit cube, the scale of the 3D model must be adjusted with respect to the input 3D hand.
@@ 5.1 Rigid Alignment
The remaining parameters to be optimized are object rotation $R$ and translation $T$. In order to (1) guarantee a globally optimal solution while (2) maintaining fixed compute time, we discretize the space of all possible parameter values to construct finite state spaces $\{\mathcal{R},\mathcal{T}\}$. The 3D rotation group, SO(3), is discretized using the unit cube subdivision strategy proposed in [33]. Translation is defined as an offset from the mean location of the MCC-HO point cloud $\mu(X)$, and the space is discretized by constructing a small voxel grid centered at $\mu(X)$. In order to avoid intractably large state space sizes (e.g., $\mathbb{R}^{3}\times\mathbb{R}^{3}$), we first optimize for rotation and then translation. For each parameter, we select the state that results in the lowest error using the sum of the following two metrics.
MCC-HO Chamfer Distance. The Chamfer Distance (CD) between the MCC-HO point cloud and a transformed object model is the usual 3D evaluation metric as defined in Equation 8. Note that when optimizing for rotation, the object model (centered at the origin) is first rotated and then translated to the mean location of the MCC-HO point cloud.
DINOv2 PCA Similarity. The MCC-HO point cloud gives us a reasonable estimate for object pose, but using this signal alone may not result in poses that align well with the input image (particularly for geometry with symmetries and in-the-wild objects). Thus, we additionally use DINOv2 [52] features to determine visual similarities between a transformed object model and the input image. For each rotation/translation state, the transformed object mesh is rendered using Pytorch3D [64]. A PCA basis is constructed using the DINOv2 features of the input image, and the first three PCA components are computed for both the rendered features $\mathcal{F}_{j}$ and input image features $\mathcal{F}_{0}$. We compute the cosine similarity between the two as
${\textstyle E_{DINO}=1-\frac{1}{2}\left(\frac{\text{PCA}(\mathcal{F}_{j})\cdot\text{PCA}(\mathcal{F}_{0})}{\max(||\text{PCA}(\mathcal{F}_{j})||_{2}\cdot||\text{PCA}(\mathcal{F}_{0})||_{2}),\epsilon)}+1\right)}$ (6)
where $\mathcal{F}_{j}$ and $\mathcal{F}_{0}$ are both masked by the input image object silhouette. When ground truth object masks are not available, we use SAM 2 [65]. See Figure 3 for rotation results using example images from 100DOH [71].
@@ 6 From Images to Videos
Given a sequence of video frames with timesteps $t$, we jointly optimize for per-frame rotations $R(t)$ and subsequently per-frame translations $T(t)$. The typical approach for temporal alignment is to use gradient-based optimization  [55, 95], which is prone to issues with local minima and has unbounded computation time. Aiming for robustness and efficiency, we propose using the Viterbi algorithm [80, 15, 29] to solve for a global solution to rigid object alignment across all frames of a video sequence. The Viterbi algorithm is a dynamic programming approach to obtaining the maximum a posteriori probability of the most likely sequence of hidden states $Q=q_{1},q_{2},\ldots,q_{N}$ that results in a sequence of observed events $O=o_{1},o_{2},\ldots,o_{N}$. The hidden states $q_{t}$ can take on values from a finite set of parameters $\mathcal{Q}$ (where in our case, $\mathcal{Q}=\mathcal{R}$ or $\mathcal{Q}=\mathcal{T}$). Following the single image case, the observations $o_{t}$ are a combination of the MCC-HO inferred object point clouds and the DINOv2 image features.
Each cell of the Viterbi trellis, $v_{t}(j)$, is the probability of ending up in state $j$ after seeing the first $t$ observations and passing through the most probable state sequence $q_{1},\ldots,q_{t-1}$. Given the probability of being in every state at time $t-1$, the Viterbi probabilities are computed via:
@@ 7 Experiments
Section 7.1 details the datasets that were used for training and evaluation. Section 7.2 defines the evaluation metrics used for comparisons. We also quantitatively (Section 7.3) and qualitatively (Section 7.4) evaluate against existing work and find that our approach achieves state-of-the-art performance on all three datasets used [4, 2, 41]. Also in Section 7.4.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach applied to in-the-wild Internet videos from 100DOH [71] to demonstrate the feasibility and scalability of our automated model retrieval and fitting. Section 7.5 includes ablation studies for MCC-HO, RAR, and each component of our Viterbi temporal alignment algorithm.
@@
F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
HO [23] 0.24 0.48 4.76 0.03 0.06 49.8 0.28 0.51 3.86
IHOI [94] - - - 0.13 0.24 23.1 0.42 0.70 2.7
HORSE [58] 0.20 0.41 7.16 0.08 0.20 27.2 0.21 0.41 7.68
MCC-HO (Ours) 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
Table 1: We compare our method, MCC-HO, to prior works on held-out test images from DexYCB, MOW, and HOI4D. Chamfer Distance (cm2) and F-score (5mm, 10mm) are reported.
@@
DiffHOI [89] 0.62 0.91 0.8
G-HOP [96] 0.76 0.97 0.4
MCC-HO + RAR (Ours) 0.91 0.98 0.28
Table 2: Quantitative evaluation of video-based hand-held object reconstruction methods on the HOI4D test dataset. Mean metrics are reported for consistency with prior work.
@@ 7.1 Datasets
HOI4D. The HOI4D dataset [41] consists of 4,000 lab videos of hands interacting with specified categories of common objects. We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets. In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [95]. For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
DexYCB. The DexYCB dataset [4] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences.
MOW/RHOV. MOW [2] is a dataset of 512 images from 100 Days of Hands [71] labeled with 3D hands and objects. We use the same splits as [94] with 350 training images and 92 test images. Additionally, we use 110 videos from RHOV [55], each of which is a 2 second clip centered at the corresponding MOW frame. In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
100DOH. The 100 Days of Hands dataset [71] contains 131 days of footage collected from Internet videos showing hands interacting with objects. The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
For DexYCB [4], MOW [2], and HOI4D [41], all results generated using our approach use the ground truth 3D hand provided by each dataset. For 100DOH [71] results (e.g., examples not contained in MOW), we use HaMeR [56] to estimate the 3D hand.
@@ 7.2 Evaluation Metrics
Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics: CD (cm2) and F-score reported at 5mm and 10mm. Both metrics are computed on point sets. The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [69] and HOI4D optimizes hand pose using 2D keypoint annotations. For each example, we sample 10k points on the surface of the ground truth mesh consistent with [94, 95] (no transformations are applied to the ground truth geometry). The MCC-HO predicted point clouds are supersampled/subsampled to obtain 10k points as well. In order to account for unknown ground truth camera parameters, ICP with scaling [64] is used to align all predicted geometry with the ground truth mesh. Median metrics are reported unless otherwise noted.
Chamfer Distance. The CD metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
$CD(A,B)=\frac{1}{N}\sum_{a\in A}\min_{b\in B}||a-b||^{2}_{2}+\frac{1}{N}\sum_{b\in B}\min_{a\in A}||a-b||^{2}_{2},$ (8)
where the point sets should have the same size. This equation is consistent with prior work [5, 6, 94, 95]. We report CD in cm2 as in[5, 6].
F-score. The F-score at a given threshold $t$ is defined as a function of precision $P$ and recall $R$:
$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{P(X_{pd},X_{gt},t)\cdot R(X_{pd},X_{gt},t)}{P(X_{pd},X_{gt},t)+R(X_{pd},X_{gt},t)}.$ (9)
@@ 7.3 Quantitative Evaluation
We quantitatively evaluate MCC-HO on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [23, 94, 58]. See Table 1. In order to sample points more densely in hand-object regions, the dimensions and granularity of the voxel grid used to query MCC-HO are determined from an initial coarse prediction using a default voxel grid. For DexYCB and HOI4D, we use every 5th frame of each test video sequence. The results for IHOI [94] and HO [23] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [95] papers; otherwise, we used available public repositories of prior work. Because AlignSDF [5] and gSDF [6] trained DexYCB-specific models, whereas we train a single model on multiple datasets, we do not compare to their published metrics. Our method achieves state-of-the-art performance for the Chamfer Distance and F-score metrics on all three datasets, implying that our technique is not overfitted to any particular dataset. We conjecture that increasing the number of unique object instances seen during training is important to model generalization. In addition, the pretrained MCC model provides a strong object prior for initialization.
We also compare to existing video-based approaches for RGB hand-held object reconstruction in Table 2, noting that the ground truth object is used for our RAR rigid alignment (a reasonable comparison as DiffHOI [95] and GHOP [96] train category-specific diffusion models using the ground truth HOI4D 3D objects).
@@ 7.4 Qualitative Evaluation
We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well retrieved object models using RAR can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks). Figure 4 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud. Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera. Because the object geometry is conditioned on the 3D hand, our approach tends to predict reasonable hand-object contact. Additionally, the object in contact with the hand is correctly identified by MCC-HO without input segmentation maps.
Figure 4: MCC-HO results for MOW test examples. The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
Figure 5: Our approach applied to images in the 100 Days of Hands dataset [71]. For each image, we estimate 3D hands using HaMeR [56], inference object geometry using MCC-HO, retrieve a 3D model using RAR (e.g., GPT-4(V) and Genie), and rigidly align the 3D model with the inferred point cloud from MCC-HO.
@@ 7.4.1 Retrieval-Augmented Reconstruction
The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images and videos. Figure 5 shows our results for 9 images from 100DOH [71] that do not have 3D ground truth (e.g., not in MOW). Estimated 3D hands for each image are obtained using HaMeR [56]. At test time, our method does not require an input hand-object segmentation mask [2, 55], depth image [89], or hand joint locations [94, 8]. Please see the supplementary materials for additional results.

### v4

@@ 5 Retrieval-Augmented Reconstruction
Like any other deep learning model, how well MCC-HO is expected to perform is related to the amount of training data available, e.g., the size of available hand-object datasets. These are currently all quite small, orders of magnitude smaller than the “big data” on which computer vision foundation models have been trained. How might we find a way to exploit these other models to our end? Specifically, recognition of objects is very well developed [gpt4v], and large collections of 3D object models [deitke2023objaverse, deitke2024objaverse] exist and have facilitated advancements in 3D generative modeling [luma, poole2022dreamfusion, chen2024text, long2024wonder3d]. Thus, in order to further improve upon network-inferred object geometry, we exploit these “big data” models by using GPT-4(V) [gpt4v] and a text-to-3D generative model [luma] to automatically detect and “retrieve” a 3D object corresponding to the hand-held object in an image. We call this technique Retrieval-Augmented Reconstruction (RAR), somewhat in analogy to Retrieval-Augmented Generation in natural language processing [lewis2020retrieval].
Given a representative frame of the video, we prompt GPT-4(V) to provide a detailed description of the object the hand is holding. This text caption is passed to Genie, a text-to-3D model from Luma [luma], to retrieve realistic 3D object geometry and appearance (Genie is also used in [li2023object]). Since Genie outputs geometry normalized to fit inside a unit cube, the scale of the 3D model must be adjusted with respect to the input 3D hand.
@@ 5.1 Rigid Alignment
The remaining parameters to be optimized are object rotation $R$ and translation $T$. In order to (1) guarantee a globally optimal solution while (2) maintaining fixed compute time, we discretize the space of all possible parameter values to construct finite state spaces $\{\mathcal{R},\mathcal{T}\}$. The 3D rotation group, SO(3), is discretized using the unit cube subdivision strategy proposed in [kurz2017discretization]. Translation is defined as an offset from the mean location of the MCC-HO point cloud $\mu(X)$, and the space is discretized by constructing a small voxel grid centered at $\mu(X)$. In order to avoid intractably large state space sizes (e.g., $\mathbb{R}^{3}\times\mathbb{R}^{3}$), we first optimize for rotation and then translation. For each parameter, we select the state that results in the lowest error using the sum of the following two metrics.
MCC-HO Chamfer Distance. The Chamfer Distance (CD) between the MCC-HO point cloud and a transformed object model is the usual 3D evaluation metric. Note that when optimizing for rotation, the object model (centered at the origin) is first rotated and then translated to the mean location of the MCC-HO point cloud.
DINOv2 PCA Similarity. The MCC-HO point cloud gives us a reasonable estimate for object pose, but using this signal alone may not result in poses that align well with the input image (particularly for geometry with symmetries and in-the-wild objects). Thus, we additionally use DINOv2 [oquab2023dinov2] features to determine visual similarities between a transformed object model and the input image. For each rotation/translation state, the transformed object mesh is rendered using Pytorch3D [ravi2020accelerating]. A PCA basis is constructed using the DINOv2 features of the input image, and the first three PCA components are computed for both the rendered features $\mathcal{F}_{j}$ and input image features $\mathcal{F}_{0}$. We compute the cosine similarity between the two as
$E_{DINO}=1-\frac{1}{2}\left(\frac{\text{PCA}(\mathcal{F}_{j})\cdot\text{PCA}(\mathcal{F}_{0})}{\max(||\text{PCA}(\mathcal{F}_{j})||_{2}\cdot||\text{PCA}(\mathcal{F}_{0})||_{2}),\epsilon)}+1\right)$} (6)
where $\mathcal{F}_{j}$ and $\mathcal{F}_{0}$ are both masked by the input image object silhouette. When ground truth object masks are not available, we use SAM 2 [ravi2024sam]. See Figure 3 for rotation results using example images from 100DOH [shan2020understanding].
@@ 6 From Images to Videos
Given a sequence of video frames with timesteps $t$, we jointly optimize for per-frame rotations $R(t)$ and subsequently per-frame translations $T(t)$. The typical approach for temporal alignment is to use gradient-based optimization  [patel2022learning, ye2023diffusion], which is prone to issues with local minima and has unbounded computation time. Aiming for robustness and efficiency, we propose using the Viterbi algorithm [viterbi1967error, forney1973viterbi, jurafsky2000speech] to solve for a global solution to rigid object alignment across all frames of a video sequence. The Viterbi algorithm is a dynamic programming approach to obtaining the maximum a posteriori probability of the most likely sequence of hidden states $Q=q_{1},q_{2},\ldots,q_{N}$ that results in a sequence of observed events $O=o_{1},o_{2},\ldots,o_{N}$. The hidden states $q_{t}$ can take on values from a finite set of parameters $\mathcal{Q}$ (where in our case, $\mathcal{Q}=\mathcal{R}$ or $\mathcal{Q}=\mathcal{T}$). Following the single image case, the observations $o_{t}$ are a combination of the MCC-HO inferred object point clouds and the DINOv2 image features.
Each cell of the Viterbi trellis, $v_{t}(j)$, is the probability of ending up in state $j$ after seeing the first $t$ observations and passing through the most probable state sequence $q_{1},\ldots,q_{t-1}$. Given the probability of being in every state at time t1t-1, the Viterbi probabilities are computed via:
@@ 7 Experiments
Section 7.1 details the datasets that were used for training and evaluation. Section 7.2 defines the evaluation metrics used for comparisons. We also quantitatively (Section 7.3) and qualitatively (Section 7.4) evaluate against existing work and find that our approach achieves state-of-the-art performance on all three datasets used [chao2021dexycb, cao2021reconstructing, liu2022hoi4d]. Also in Section 7.4.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach applied to in-the-wild Internet videos from 100DOH [shan2020understanding] to demonstrate the feasibility and scalability of our automated model retrieval and fitting. Section 7.5 includes ablation studies for MCC-HO, improvements attributed to RAR, and the effect of conditioning hand-held object reconstruction on 3D hands.
@@
F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) Vol ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) Vol ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) Vol ($\downarrow$)
HO [hasson2019learning] 0.24 0.48 4.76 11.8 0.03 0.06 49.8 78.2 0.28 0.51 3.86 12.7
IHOI [ye2022s] - - - - 0.13 0.24 23.1 20.4 0.42 0.70 2.7 0.91
HORSE [Prakash2024HOI] 0.23 0.42 6.97 3.63 0.11 0.23 24.5 17.0 0.26 0.45 6.69 1.39
MCC-HO (Ours) 0.36 0.60 3.74 2.42 0.15 0.31 15.2 13.8 0.52 0.78 1.36 1.94
Table 1: We compare our method, MCC-HO, to prior works on held-out test images from DexYCB, MOW, and HOI4D. Chamfer Distance (cm2), F-score (5mm, 10mm), and Intersection Volume using 0.5cm voxels (as in HO [hasson2019learning], IHOI [ye2022s]) are reported.
@@
G-HOP [ye2024g] 0.61 0.89 0.8
MCC-HO + RAR 0.74 0.91 0.64
Table 2: Quantitative comparison to a state-of-the-art video-based hand-held object reconstruction method using the HOI4D test dataset. Mean metrics are reported for consistency with G-HOP.
@@ 7.1 Datasets
HOI4D. The HOI4D dataset [liu2022hoi4d] consists of 4,000 lab videos of hands interacting with specified categories of common objects. We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets. In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [ye2023diffusion]. For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
DexYCB. The DexYCB dataset [chao2021dexycb] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences.
MOW/RHOV. MOW [cao2021reconstructing] is a dataset of 512 images from 100 Days of Hands [shan2020understanding] labeled with 3D hands and objects. We use the same splits as [ye2022s] with 350 training images and 92 test images. Additionally, we use 110 videos from RHOV [patel2022learning], each of which is a 2 second clip centered at the corresponding MOW frame. In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
100DOH. The 100 Days of Hands dataset [shan2020understanding] contains 131 days of footage collected from Internet videos showing hands interacting with objects. The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
For DexYCB [chao2021dexycb], MOW [cao2021reconstructing], and HOI4D [liu2022hoi4d], all results generated using our approach use the ground truth 3D hand provided by each dataset. For 100DOH [shan2020understanding] results (e.g., examples not contained in MOW), we use HaMeR [pavlakos2023reconstructing] to estimate the 3D hand.
@@ 7.2 Evaluation Metrics
Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics: CD (cm2) and F-score reported at 5mm and 10mm. Both metrics are computed on point sets. The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [rong2020frankmocap] and HOI4D optimizes hand pose using 2D keypoint annotations. For each example, we sample 10k points on the surface of the ground truth mesh consistent with [ye2022s, ye2023diffusion] (no transformations are applied to the ground truth geometry). The MCC-HO predicted point clouds are supersampled/subsampled to obtain 10k points as well. In order to account for unknown ground truth camera parameters, ICP with scaling [ravi2020accelerating] is used to align all predicted geometry with the ground truth mesh. Median metrics are reported unless otherwise noted.
@@ 7.3 Quantitative Evaluation
We quantitatively evaluate MCC-HO on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [hasson2019learning, ye2022s, Prakash2024HOI]. See Table 1. In order to sample points more densely in hand-object regions, the dimensions and granularity of the voxel grid used to query MCC-HO are determined from an initial coarse prediction using a default voxel grid. For DexYCB and HOI4D, we use every 5th frame of each test video sequence. The results for IHOI [ye2022s] and HO [hasson2019learning] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [ye2023diffusion] papers; otherwise, we used available public repositories of prior work. Our method achieves state-of-the-art performance for the Chamfer Distance and F-score metrics on all three datasets, implying that our technique is not overfitted to any particular dataset. We conjecture that increasing the number of unique object instances seen during training is important to model generalization. Thus, the pretrained MCC model provides a strong object prior for initialization.
We also compare to G-HOP [ye2024g], a state-of-the-art video-based approach for RGB hand-held object reconstruction in Table 2. Our method (using generated Genie objects for RAR) outperforms the generalist G-HOP model (that was trained on seven hand-object datasets [brahmbhatt2020contactpose, chao2021dexycb, corona2020ganhand, liu2022hoi4d, taheri2020grab, yang2022oakink]) in all three metrics.
@@ 7.4 Qualitative Evaluation
We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well retrieved object models using RAR can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks). Figure 4 shows predicted point clouds for a number of MOW test examples, as well as the result of RAR. Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera. Because the object geometry is conditioned on the 3D hand, our approach tends to predict reasonable hand-object contact. Additionally, the object in contact with the hand is correctly identified by MCC-HO without input segmentation maps. Figure 5 shows qualitative comparisons to existing work, where the combination of MCC-HO + RAR allows us to obtain more realistic object geometry.
Figure 4: MCC-HO results for MOW test examples. The input image (top, left), network-inferred hand-object point cloud (top, right), RAR (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
Figure 5: Qualitative comparisons using MOW test examples. *Note that the MCC-HO point clouds are rendered as a mesh via Poisson surface reconstruction, which can introduce artifacts not attributed to our method. The last column is MCC-HO + RAR.
Figure 6: Our approach applied to in-the-wild images from the 100 Days of Hands dataset [shan2020understanding]. For each image, we estimate 3D hands using HaMeR [pavlakos2023reconstructing], inference object geometry using MCC-HO and retrieve a 3D model using RAR (e.g., GPT-4(V) and Genie).
@@ 7.4.1 Retrieval-Augmented Reconstruction
The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images and videos. Figure 6 shows our results for 9 images from 100DOH [shan2020understanding] that do not have 3D ground truth (e.g., not in MOW). Estimated 3D hands for each image are obtained using HaMeR [pavlakos2023reconstructing]. At test time, our method does not require an input hand-object segmentation mask [cao2021reconstructing, patel2022learning], depth image [wu2023multiview], or hand joint locations [ye2022s, choi2023handnerf]. Please see the supplementary materials for additional results.